### v1

# Can LLMs Simulate Human Behavioral Variability? A Case Study in the Phonemic Fluency Task

Abstract
Large language models (LLMs) are increasingly explored as substitutes for human participants in cognitive tasks, but their ability to simulate human behavioral variability remains unclear. This study examines whether LLMs can approximate individual differences in the phonemic fluency task, where participants generate words beginning with a target letter. We evaluated 34 model configurations, varying prompt specificity, sampling temperature, and model type, and compared outputs to responses from 106 human participants. While some configurations, especially Claude 3.7 Sonnet, matched human averages and lexical preferences, none reproduced the scope of human variability. LLM outputs were consistently less diverse and structurally rigid, and LLM ensembles failed to increase diversity. Network analyses further revealed fundamental differences in retrieval structure between humans and models. These results highlight key limitations in using LLMs to simulate human cognition and behavior.
Can LLMs Simulate Human Behavioral Variability? A Case Study in the Phonemic Fluency Task
Mengyang Qiu1††thanks: Corresponding author. Zoe Brisebois1 Siena Sun2 1Department of Psychology, Trent University, Canada 2Speech-Language Pathology Program, Saint Elizabeth University, United States {mengyangqiu, zoebrisebois}@trentu.ca ssun@steu.edu
1 Introduction
Large language models (LLMs) have rapidly advanced in recent years, achieving impressive performance across a wide range of natural language tasks. As a result, researchers have become increasingly interested in using LLMs as experimental tools in cognitive and behavioral science. Some even propose that LLMs could replace human participants in certain studies, offering scalable and efficient alternatives for simulating human behavior (Dillion et al., 2023).
This idea has intuitive appeal: LLMs can generate fluent responses on demand, and arguably encode the “wisdom of the crowd” from their massive training data (Trott, 2024). For instance, Hansen and Hebart (2022) found that GPT-3 could generate semantic features for concepts that not only mirrored the distribution of human-generated features but also matched human norms in their ability to predict similarity, relatedness, and category membership. On the other hand, a growing body of work has pointed out that LLMs may lack a crucial feature of human language and cognition: variability. Cuskley et al. (2024) argue that LLMs are limited by their training on written language, which captures only a narrow slice of human communicative behavior. Zanotto and Aroyehun (2024) offer empirical support, showing that LLM-generated texts exhibit substantially lower linguistic diversity than human-written texts across a range of features, including syntax, vocabulary, and style.
This lack of variability also becomes evident when examining how LLMs perform in tasks designed to probe semantic or associative structures. In a semantic fluency task where participants named as many animals as possible, Wang et al. (2025) found that LLM-generated semantic networks were structurally different from those of humans. Compared to human networks, those of LLMs exhibited weaker local associations, poorer global integration, and greater rigidity in semantic organization. Similarly, Haim et al. (2025) examined word association networks in a STEM-related mindset task and found that GPT-3.5 produced networks that were notably sparser and less interconnected than those generated by human participants.
It is worth noting that Wang et al. (2025) recognized this limitation and attempted to address it by prompting LLMs to role-play 30 different occupations, effectively simulating multiple distinct agents. However, the resulting semantic networks still failed to match the flexibility and associative structure of human data. Crucially, real human participants, even those drawn from relatively homogeneous groups such as college students, consistently exhibit meaningful individual differences. These differences may reflect not only variation in the amount, type, and organization of semantic knowledge, but also differences in executive function, and strategy use. Of course, because LLMs are optimized to encode and reproduce stable semantic relationships, consistency across prompts may reflect a design feature. From this perspective, the internal coherence and semantic regularity that make LLMs so effective at language modeling may also constrain their ability to simulate the variability that characterizes human cognition.
If semantic-based tasks limit LLMs’ ability to simulate human behavioral variability, it raises the question: how would they perform on a task that relies less on meaning? One such task is phonemic (or letter) fluency, where individuals are asked to produce as many words as possible that begin with a particular letter, such as F, within a fixed time limit. Unlike semantic fluency, this task requires an effortful search through the mental lexicon based on orthographic or phonological form–an unnatural retrieval strategy in everyday language use. Prior analyses have shown that phonemic fluency performance is influenced by both lexical properties, such as word frequency, age of acquisition, and phonetic similarity, and cognitive factors like switching patterns (Cho et al., 2021).
The goal of the present study is to explore whether LLMs are more capable of simulating inter-participant variability in a less semantically driven challenge–namely, the phonemic fluency task–or whether their outputs remain rigid, perhaps driven by frequency-based or alphabetic heuristics. Alongside this primary question, we examine several additional factors that may influence the diversity and human-likeness of LLM-generated responses:
1.
Participant context prompts: We systematically vary the type of information provided to the model, including demographic details (age, education) and performance-level information (number of correct responses), to assess whether real participant metadata supports more individualized outputs.
2.
Sampling temperature: Building on prior work that links temperature to creative or divergent generation (Peeperkorn et al., 2024), we test whether adjusting this parameter affects response diversity and whether higher temperatures introduce human-like variation or merely add noise.
3.
Model architecture and reasoning capacity: We compare general-purpose LLMs (e.g., GPT-4.1) to reasoning-oriented models (e.g., O3-mini) to examine whether models designed to engage in multi-step thinking produce more flexible or human-aligned behavior.
4.
Ensemble production: Finally, we investigate whether sampling outputs from multiple models, rather than relying on single model outputs, can better approximate the range of responses observed across human participants. While recent ensemble methods have primarily focused on improving task accuracy by using techniques like sampling and majority voting to converge on a correct answer (Li et al., 2024), our goal is the opposite: to explore whether ensembling can simulate the distributional variability seen in human behavior.
In preview, our findings suggest a clear answer: LLMs are highly capable of producing fluent and correct answers in the phonemic fluency task, but none of them match the full scope of human behavioral variability. The remainder of the paper is organized as follows. Section 2 describes the human F fluency dataset, the full range of LLM configurations tested, and our approach to simulating participant-level outputs under different prompt and temperature conditions. Section 3 evaluates how well LLMs replicate human response counts and lexical diversity across different prompting setups and models. Section 4 presents item-level analyses, examining how linguistic features (e.g., word frequency, age of acquisition, phonological neighbors) predict production frequency in humans versus models. Section 5 turns to structural comparisons of word co-occurrence networks, analyzing clustering, path length, and representational similarity. Section 6 concludes with a discussion of implications for using LLMs as behavioral proxies and offers recommendations for improving variability in future simulations.
2 Method
2.1 Human data source
We used verbal fluency data collected by Qiu and Johns (2021)111The verbal fluency data are publicly available on the Open Science Framework under the Creative Commons Attribution 4.0 International License (CC BY 4.0). We adhered to all applicable terms for reuse and attribution., who investigated noun- and verb-based semantic fluency across two experiments. In both experiments, participants first completed a phonemic fluency task using the letter F as a familiarization trial before proceeding to semantic tasks. A total of 106 native English speakers (age: ; education: years, ) were recruited via Amazon Mechanical Turk and completed the tasks online. Audio recordings were obtained for all participants’ responses to the F fluency task.
Following the original study’s procedure, we used the Google Speech-to-Text API222https://cloud.google.com/speech-to-text to transcribe the recordings. Transcripts were then manually reviewed to identify and exclude errors, including repetitions, out-of-category responses, same-lemma variants (e.g., fish, fishing, fished), and counting behavior, where multiple number words (e.g., four, five, fourteen) were produced.
Participants produced an average of 17.27 words (), which slightly decreased to 16.89 words () after manual corrections. This small reduction indicates overall good adherence to task instructions. However, this average was lower than participants’ performance on the semantic fluency tasks in the same study, especially the noun-based conditions (animals, and living or nonliving things), where the mean number of responses exceeded 22. This pattern highlights the greater difficulty of phonemic fluency relative to semantic fluency.
2.2 LLM simulation procedure
We evaluated a diverse set of LLMs from six major providers. For closed-source models, we used the official APIs from OpenAI333https://openai.com/api/, Anthropic444https://www.anthropic.com/api, Google555https://ai.google.dev/gemini-api/docs, and xAI666https://x.ai/api. For open-source models, we accessed models from Meta and Alibaba via the Together.ai API777https://www.together.ai. Each configuration was prompted to simulate all 106 human participants.888All LLM-generated outputs used in this study will be made publicly available under the Creative Commons Attribution 4.0 International License (CC BY 4.0).
To assess behavioral realism, we systematically varied prompt specificity and sampling temperature. A representative full prompt is shown in Figure 1. This version included the participant age, education, and number of correct responses. We also tested three reduced-information variants: a demographic-only prompt (no number of correct responses), a performance-only prompt (no demographics), and a no-information prompt. All prompts followed the same task format and instruction wording as the human study, and models were instructed to output one word per line.
In total, 34 distinct model configurations were evaluated. Appendix A provides detailed specifications for each model, including provider, temperature settings, and prompt type.
3 Participant-Level Analysis
3.1 Number of responses
Models from leading LLM providers were evaluated on their ability to simulate human performance in the letter F fluency task, with additional testing of the latest models from OpenAI (GPT-4.1) and Anthropic (Claude 3.7 Sonnet) using limited participant information. Human participants produced an average of 16.89 correct responses within the one-minute time constraint (). Most model configurations, when provided with full information or at least performance information (i.e., number of correct responses), generated similar response counts to humans (after removing repetitions and other errors), suggesting generally successful adherence to the instructed constraints.
When provided with only demographic information (age and education), GPT-4.1 generated significantly more words () and Claude 3.7 Sonnet showed similar overproduction ()–both more than doubling human output. This pattern was even more pronounced in the no-information condition, with GPT-4.1 generating higher counts () and Claude 3.7 Sonnet producing comparable results ().
Surprisingly, O4-mini, one of OpenAI’s latest reasoning models, demonstrated extreme overestimation with demographic information ()–nearly 60 more words than human participants. Even when performance information was explicitly provided to this model, it still overproduced with considerable variance (), indicating inconsistent adherence to the performance constraints specified in the prompt. Interestingly, with the full prompt that included both demographic and performance information, O4-mini accurately reproduced human-like performance in terms of response count.
For several other models, even with full participant information, we observed unstable production patterns characterized by extreme bursts of responses, sometimes generating around 100 words for some participants while appropriately following instructions for others. In particular, OpenAI’s GPT-4-turbo () and O3 () models generated excessive output with high variance, suggesting these models inconsistently applied the performance constraints specified in the prompt.
Table 1 summarizes the 11 model configurations that significantly overestimated human performance at the group level. Additionally, four more configurations, including Google’s Gemini 2.0 and Gemini 2.5 (with and without thinking enabled), and Alibaba’s Qwen-2.5-72B, showed similar bursts of responses for individual participants, though these did not reach statistical significance at the group level. These 15 model configurations were excluded from subsequent analyses to focus on models with more realistic human simulation.
Provider | Model | Prompt Type | Mean SD | Diff. | t |
Human | – | – | – | – | |
Models with Demographic-Info or No-Info Prompts | |||||
OpenAI | GPT-4.1 | Demographic | |||
No | |||||
O4-mini (reasoning) | Demographic | ||||
[…]
No | |||||
Models with Performance-Info or Full Prompts | |||||
OpenAI | GPT-4-turbo | Full | |||
O3 (reasoning) | Full | ||||
O4-mini (reasoning) | Performance | ||||
[…]
Llama 3.3 70B | Full | ||||
Note: Diff. = mean difference from human performance; ; |
3.2 Variability of responses
To evaluate the variability of responses across human participants and LLM-simulated participants, we computed the Type-to-Token Ratio (TTR) and the Idiosyncratic Type-to-Total Type Ratio (ITTR). TTR captures lexical diversity by dividing the number of unique words (types) by the total number of responses (tokens). A higher TTR indicates a wider range of vocabulary produced. ITTR measures how many of those types were idiosyncratic (i.e., produced by only one participant) relative to the total number of types. This provides insight into how consistent or individualized the generated words are across participants within the group (Castro et al., 2021).
As shown in Table 2, human participants produced the highest number of unique types (476) and idiosyncratic types (201), resulting in a TTR of 0.27 and an ITTR of 0.42. No LLMs approached this level of variability. The most diverse LLM output came from Claude 3.7 Sonnet at temperature 1.0 (226 types), followed closely by O3-mini, the best-performing reasoning model (184 types). However, their ITTRs (0.32 and 0.28, respectively) remained well below the human benchmark.
As expected, increasing temperature led to more diverse responses for both GPT-4.1 and Claude 3.7 Sonnet, with higher TTR and ITTR values at higher temperatures, though they still failed to match human-like variability. We also tested how prompt specificity affected variability: for GPT-4.1, removing demographic details while retaining performance information slightly increased TTR and ITTR, but for Claude 3.7, the pattern reversed. A trend across model generations also emerged: holding temperature constant at 0.7, lexical diversity decreased from GPT-3.5 to GPT-4o to GPT-4.1. This suggests that newer models may prioritize consistency and safety over variability.
Finally, we examined whether ensembles of LLMs could recover some of the lost variability by combining outputs from diverse models. Three ensemble strategies were tested: sampling from the top five models in terms of type count, from all models with at least 100 types (9 in total), and from all models that received the full prompt (17 in total). As shown in the bottom portion of Table 2, none of the ensemble approaches exceeded the lexical diversity of the best individual model. The top-five ensemble achieved 225 types, nearly identical to Claude 3.7 Sonnet at temperature 1.0. This suggests that between-model differences in lexical choices are limited, and ensemble generation alone is insufficient to reach human-level diversity.
Provider | Model | Temperature | Types | Idio. Types | Tokens | TTR | ITTR |
Human | – | 476 | 201 | 1790 | 0.27 | 0.42 | |
OpenAI | |||||||
GPT-4.1 | 0.3 | 57 | 10 | 1766 | 0.03 | 0.18 | |
0.7 (default) | 69 | 18 | 1778 | 0.04 | 0.26 | ||
1.1 | 108 | 37 | 1783 | 0.06 | 0.34 | ||
1.5 | 138 | 38 | 1777 | 0.08 | 0.28 | ||
[…]
GPT-4o | 0.7 (default) | 128 | 38 | 1773 | 0.07 | 0.30 | |
GPT-3.5 | 0.7 (default) | 148 | 48 | 1783 | 0.08 | 0.32 | |
[…]
O4-mini (reasoning) | – | 88 | 24 | 1790 | 0.05 | 0.27 | |
Anthropic | |||||||
Claude 3.7 Sonnet | 0.3 | 97 | 28 | 1787 | 0.05 | 0.29 | |
0.7 | 150 | 39 | 1788 | 0.08 | 0.26 | ||
1.0 | 226 | 73 | 1781 | 0.13 | 0.32 | ||
[…]
Claude 3.7 Sonnet
0.7 | 137 | 35 | 1787 | 0.08 | 0.26 |
Claude 3.7 Sonnet
(thinking enabled) |
– | 137 | 43 | 1788 | 0.08 | 0.31 | |
xAI | |||||||
Grok-3 | 0.7 | 100 | 22 | 1790 | 0.06 | 0.22 | |
1.0 | 94 | 18 | 1790 | 0.05 | 0.19 | ||
[…]
1.0 | 71 | 19 | 1793 | 0.04 | 0.27 | ||
Open-Source (Meta) | |||||||
Llama 4 Scout | default | 83 | 25 | 1783 | 0.05 | 0.30 | |
LLM Ensembles | |||||||
Top 5 Models
(Word Types) |
– | 225 | 75 | 1778 | 0.13 | 0.33 | |
100+ Word Types
(9 models) |
– | 180 | 46 | 1785 | 0.10 | 0.26 | |
All-Model Mix
(17 models) |
– | 169 | 45 | 1780 | 0.09 | 0.27 |
4 Item-Level Analysis
4.1 Distribution of production frequency
Previous studies have observed that word production in verbal fluency tasks typically follows Zipf’s law, a type of power-law relationship where word frequency is inversely proportional to its rank (e.g., Taler et al., 2020). This distribution is characterized by a small number of words that are produced frequently across participants, followed by a “long tail” of words that are produced infrequently, often appearing only once or twice in the entire dataset. In the current study, human participants’ word production closely followed this expected pattern, with an coefficient of and a strong goodness of fit (). The most frequently produced word by human participants was “fun”, mentioned 34 times, followed by “fire” (30 times) and “fan” (28 times).
In contrast, LLMs, both single and ensembled models, demonstrated a notably different pattern. While all models showed good fit to a power-law distribution ( ranging from to ), their higher coefficients (–) reflect a steeper decline in frequency, with top-ranked words dominating the distribution more heavily than in human data. For instance, O3-mini produced “fast” 83 times, and All-Model Mix generated “fish” 90 times–frequencies substantially higher than those observed in human responses. The Top 5 Models ensemble demonstrated the closest fit to the human distribution patterns (), though still with a steeper coefficient (). Table 3 presents the Zipf’s law goodness-of-fit statistics for all models.
Model | ||
Human | 0.89 | 0.92 |
Claude 3.7 Sonnet | 1.19 | 0.88 |
O3-mini | 1.25 | 0.87 |
[…]
Models with 100+ Word Types | 1.22 | 0.88 |
All-Model Mix | 1.28 | 0.90 |
4.2 Linguistic variables and production frequency
Previous research has shown that verbal fluency performance is influenced by multiple linguistic variables. For instance, Taler et al. (2020) reported that semantic neighborhood size and word frequency were significant predictors of word production in semantic fluency tasks, with words from denser semantic neighborhoods and higher frequency being produced more often. Similarly, Cho et al. (2021) found that phonemic fluency production was influenced by lexical-semantic factors such as word frequency, familiarity, word duration, and age of acquisition.
In this study, we examined the relationship between production frequency and core linguistic variables obtained from the English Lexicon Project (Balota et al., 2007), including word length (number of letters), word frequency (log frequency per million words from SUBTLEX corpus), orthographic neighborhood size (number of words that differ by one letter), phonological neighborhood size (number of words that differ by one phoneme), semantic neighborhood size (number of semantically related words), and age of acquisition (estimated age when a word is typically learned). We first conducted simple correlation analyses between these variables and production frequency.
As shown in Table 4, all linguistic variables except for semantic neighborhood size were significantly correlated with frequency of responses in human data. Word frequency showed the strongest positive correlation (), indicating that high-frequency words were more likely to be produced. Word length and age of acquisition were negatively correlated with production frequency ( and , respectively), suggesting a preference for shorter and earlier-acquired words. Both orthographic and phonological neighborhood sizes, often closely correlated and sometimes used interchangeably in psycholinguistic studies, showed positive correlations, with denser neighborhoods associated with higher production frequency.
Among the LLM models, Claude 3.7 Sonnet most closely approximated the human pattern, with highly similar correlation coefficients across variables. O3-mini followed a similar trend but exhibited generally weaker correlations, particularly for orthographic neighborhood size. The ensemble models followed the same general trend but with progressively weaker correlations, especially for the phonological and orthographic neighborhood variables. Despite differences in correlation strength, the ranking of variable importance was consistent across models: word frequency and age of acquisition showed the strongest correlations, while semantic neighborhood size remained non-significant. This aligns with the phonemic nature of the task, which prioritizes phonological over semantic retrieval processes–a contrast to semantic fluency tasks where semantic factors are more predictive (Taler et al., 2020).
Model | Length | WF | Ortho_N | Phono_N | Sem_N | AoA |
Human | ∗∗∗ | ∗∗∗ | ∗∗∗ | ∗∗ | ∗∗∗ | |
Claude 3.7 Sonnet | ∗∗∗ | ∗∗∗ | ∗∗∗ | ∗∗∗ | ∗∗∗ | |
O3-mini | ∗∗∗ | ∗∗∗ | ∗∗ | ∗∗∗ | ∗∗∗ | |
[…]
Models with 100+ Word Types | ∗∗∗ | ∗∗∗ | ∗ | ∗ | ∗∗∗ | |
All-Model Mix | ∗∗∗ | ∗∗∗ | ∗∗∗ | |||
Note. Length = word length; WF = log word frequency (SUBTLEX); Ortho_N = orthographic neighborhood size; Phono_N = phonological neighborhood size; Sem_N = semantic neighborhood size; AoA = age of acquisition. ; ; |
To formally establish the factors that influence production frequency, we conducted stepwise regression analyses with all linguistic variables examined in the correlation analyses above. Full statistical details are provided in Appendix B, but a consistent pattern emerged: linguistic variables explained the most variance in production frequency for human responses (adjusted ), followed closely by Claude 3.7 Sonnet (adjusted ), with substantially less variance explained for O3-mini (adjusted ) and the ensemble models (adjusted –).
Claude 3.7 Sonnet most closely matched the human pattern in both variance explained and predictors retained, with word frequency and age of acquisition emerging as significant in both. Orthographic neighborhood size was uniquely important for humans, while word length–absent from the human model–played a stronger role in Claude and O3-mini. The ensemble models diverged more sharply, with age of acquisition emerging as the dominant–or only–predictor, and word frequency playing a minimal or negligible role. These patterns suggest both shared and divergent influences on word production across systems, with greater unpredictability in the outputs of O3-mini and ensemble models compared to humans and Claude.
5 Network Analysis
Building on our earlier findings, we focused the network analysis exclusively on comparing human and Claude 3.7 Sonnet outputs. Claude consistently exhibited the most human-like patterns–response count, variability, production frequency distribution, and linguistic predictors of word choice. In contrast, O3-mini and the ensemble models showed substantially divergent patterns with lower explained variance and different predictor profiles, suggesting fundamentally different underlying mechanisms driving their word production.
5.1 Network construction approach
To better understand the organizational structure of word production in letter F fluency, we constructed correlation-based networks–a common approach in cognitive science for examining the relationship between concepts (Borodkin et al., 2016; Li and Qiu, in press; Siew and Guru, 2023). Given the different number of nodes in the full networks (476 words for humans, 226 for Claude), direct comparison of network metrics may be confounded by size differences (Borodkin et al., 2016). Therefore, we selected 197 words that were produced by both humans and Claude participants to ensure a fairer comparison of network structure.
The network construction process involved three main steps: (1) creating binary matrices indicating which words each participant produced, (2) calculating word-word cosine similarity matrices based on co-occurrence patterns across participants, and (3) applying a thresholding procedure (TMFG; Massara et al., 2017) to identify meaningful connections between words.
5.2 Structural comparison of human and Claude networks
We compared the structural properties of networks constructed from human and Claude 3.7 Sonnet outputs using two standard global metrics: clustering coefficient (CC) and average shortest path length (ASPL). CC quantifies the degree to which nodes in a network tend to cluster together (i.e., the likelihood that two neighbors of a node are also connected), while ASPL reflects the average number of steps required to connect any two nodes, indexing the global efficiency of the network (Siew et al., 2019).
Both networks were constructed from the same set of 197 shared words to control for network size. The human network showed a higher CC () and longer ASPL () compared to Claude (, ). To assess whether the observed structures differed from random expectations, we simulated 1,000 Erdős-Rényi random networks Erdős and Rényi (1960) matched in size and density. One-sample -tests confirmed that both human and Claude networks significantly deviated from randomness, exhibiting substantially higher CCs and slightly higher ASPLs than random networks (human: for CC, for ASPL; Claude: for CC, for ASPL; all ).
To statistically compare the two networks, we conducted a subnetwork bootstrap procedure: 1,000 partial networks were generated for each system by randomly sampling 50% of nodes and recalculating network metrics (Borodkin et al., 2016; Qiu et al., 2021). As shown in Figure 2, human subnetworks had significantly higher CCs (, ) and longer ASPLs (, ) than Claude subnetworks. These differences indicate that human responses are organized into tighter local clusters with weaker global integration, whereas Claude responses form a more evenly connected network with greater global efficiency.
To complement the network analysis, we conducted a representational similarity analysis (RSA; Nili et al., 2014) comparing the original pairwise similarity matrices that underlie network construction. The Spearman correlation between the lower triangular portions of the human and Claude matrices was modest but significant (, ), providing additional evidence that Claude organizes or retrieves words differently than humans. Taken together with the network metric results, this suggests that while Claude can produce human-like words and respond to similar lexical variables, the global structure and retrieval dynamics it exhibits remain meaningfully distinct from human performance.
6 Discussion
This study evaluated whether LLMs can simulate the behavioral variability observed in human performance on a classic cognitive task: phonemic fluency. Across 34 configurations of state-of-the-art LLMs, we systematically examined the impact of prompt specificity, sampling temperature, model type, and ensemble strategies, comparing model-generated outputs to human F fluency data across three levels of analysis. While some configurations, most notably Claude 3.7 Sonnet at high temperature when prompted with full participant information, produced outputs that broadly resembled human responses in terms of average word count and sensitivity to linguistic features, no model fully captured the scope of individual differences or the structural properties characteristic of human lexical retrieval.
At the participant level, most LLM configurations achieved plausible response counts, particularly when given performance information. However, lexical diversity remained consistently lower than in human participants, and this gap persisted across temperature settings and model types. Models like Claude 3.7 Sonnet produced slightly more diverse responses than others, but even the best-performing configurations failed to reach human-level variation in word types. These findings highlight a persistent tendency toward output rigidity: models tend to repeat the same frequent words across different simulated “participants”.
This consistency persisted even when models were prompted with detailed demographic and performance information, and even when responses were sampled at higher temperatures. The role-play strategy used in prior work (Wang et al., 2025) was conceptually similar to our demographic conditioning, but both approaches appear insufficient for inducing human-like variability in model output. Moreover, ensemble outputs across LLMs, as implemented here, did not recover the distributional diversity observed in human responses. This is likely because most model configurations produced overlapping sets of frequent words. In fact, model ensembles amplified high-frequency core vocabulary and suppressed rare or idiosyncratic responses, further narrowing rather than expanding the response space.
At the item level, further differences emerged in the distribution of word production frequencies. While both human and LLM data followed power-law distributions, the steeper decline in LLM data revealed a stronger bias toward high-frequency words, with less diversity in the “long tail” of infrequent responses. When examining the linguistic factors driving word production, we found some similarities between humans and LLMs, with word frequency and age of acquisition emerging as significant predictors across systems. However, important differences were also evident. For human participants, orthographic neighborhood size played a unique and substantial role, with words having more neighbors being produced more frequently. In contrast, this factor was less influential in LLM outputs, which instead showed greater sensitivity to word length. These patterns suggest that humans may rely more heavily on form-based associations during word retrieval, while LLMs appear to prioritize more surface-level features.
Perhaps most revealing were the network analyses, which uncovered fundamental differences in how humans and LLMs organize lexical retrieval. Human responses exhibited stronger local clustering but weaker global integration, in contrast to Claude 3.7 Sonnet, the most human-like model, which optimized for network-wide efficiency. These structural mismatches underscore a core limitation of current LLMs: they can produce plausible words but fail to approximate the associative dynamics of human memory search.
Taken together, our results have important implications for researchers considering LLMs as substitutes for human participants. While LLMs excel at producing grammatical, contextually appropriate outputs, they do not capture the full spectrum of variability seen in human behavioral data. This limitation is particularly relevant for cognitive tasks where individual differences are of theoretical interest. Researchers who use LLMs as proxies for human behavior should be aware that they may significantly underestimate the true variability in their target population. That said, among the models tested, Claude 3.7 Sonnet consistently emerged as the most human-like, and performed far better than other configurations. Future work could focus on extending this model’s capabilities through more targeted persona modeling, potentially incorporating linguistic and cognitive profiles (e.g., selective reading exposure, task history) to better simulate participant-level differences.
In sum, while LLMs continue to impress in terms of fluency and surface-level plausibility, our findings highlight the challenges of using them as substitutes for human participants in behavioral research–especially when individual variability and cognitive process modeling are central. Phonemic fluency reveals these limitations clearly, offering a strong case for developing new strategies if LLMs are to be used meaningfully in the simulation of human behavior.
Limitations
This study focuses on a single verbal fluency task (letter F) in English, limiting generalizability across task types, letters, and languages. While phonemic fluency provides a stringent test of form-based retrieval, different letters may elicit distinct lexical or structural patterns. Similarly, results may differ in languages with different orthographic or phonological systems. Our analysis also centers on a specific subset of models and parameter settings; other prompting techniques (e.g., few-shot examples, chain-of-thought reasoning) or fine-tuned models may yield different patterns. Additionally, while we simulated participant-level responses using demographic and performance metadata, we did not incorporate richer individual-level traits (e.g., reading history, cognitive profiles) that may be necessary to drive human-like variability. Future work should expand to multilingual and cross-task comparisons, explore more cognitively grounded persona modeling, and evaluate whether newer architectures or sampling strategies better capture human behavioral diversity.
Appendix A Overview of LLM Providers, Prompts, and Sampling Temperatures
We evaluated 34 model configurations across six providers–OpenAI, Anthropic, Google, xAI, Meta (LLaMA), and Alibaba (Qwen)–systematically varying model family (including general-purpose vs. reasoning-focused models and different model generations), temperature, and prompt specificity. In subsequent analyses, we also examined ensemble combinations of these models to assess response variability. Table A1 summarizes these configurations. Prompt specificity ranged from full participant profiles to minimal or absent information, designed to test how different inputs affected response count and variability.
Provider | Model | Temperature | Prompt Type |
OpenAI (15 model configurations) | |||
gpt-4.1-2025-04-14 | 0.3, 0.7 (default), 1.1, 1.5 | Full (all temperatures); No Info, Demographic, Performance (0.7 only) | |
gpt-4o-2024-08-06 | 0.7 (default) | Full | |
gpt-4-turbo-2024-04-09 | 0.7 (default) | Full | |
gpt-3.5-turbo-0125 | 0.7 (default) | Full | |
[…]
o3-2025-04-16 (reasoning) | – | Full | |
o3-mini-2025-01-31 (reasoning) | – | Full | |
Anthropic (8 model configurations) | |||
claude-3-7-sonnet-20250219 | 0.3, 0.7, 1.0 (default) | Full (all temperatures); No Info, Demographic, Performance (0.7 only) | |
claude-3-5-sonnet-20241022 | 1.0 (default) | Full | |
[…]
Google (3 model configurations) | |||
gemini-2.5-flash-preview-04-17 | default | Full | |
gemini-2.0-flash | default | Full | |
[…]
xAI (4 model configurations) | |||
grok-3-latest | 0.7, 1.0 | Full | |
[…]
Open-Source Models (using together.ai API; 4 model configurations) | |||
meta-llama/Llama-4-Maverick-17B-128E-Instruct-FP8 | default | Full | |
meta-llama/Llama-4-Scout-17B-16E-Instruct | default | Full | |
meta-llama/Llama-3.3-70B-Instruct-Turbo | default | Full | |
[…]
Prompt types: Full = complete demographic and performance information; Demographic = only demographic information (age, education); No-Info = no participant information; Performance = only performance information.
Appendix B Stepwise Regression Analyses of Linguistic Predictors of Production Frequency
For human data, the stepwise regression yielded a significant model [] that accounted for of the variance (adjusted ). Word frequency emerged as the strongest predictor (), followed by orthographic neighborhood size (), and age of acquisition (). This indicates that more frequent, orthographically dense, and earlier-acquired words were produced more often in the task.
For Claude 3.7 Sonnet, stepwise regression likewise produced a significant model [] explaining of the variance in production frequency (adjusted ). As with humans, word frequency was the strongest predictor (). However, unlike humans, orthographic neighborhood size was not retained in the model; instead, word length emerged as a significant predictor (). Age of acquisition remained a significant predictor (), consistent with the human pattern. O3-mini’s regression model explained substantially less variance, only 18.8% (adjusted ) [], retaining just two predictors: word frequency () and word length ().
Among the ensemble models, the Top 5 Models ensemble yielded a significant model [] explaining of variance (adjusted ), with age of acquisition as the primary predictor () followed by word frequency (). The 100+ Word Types ensemble produced a one-predictor model [] with of variance explained (adjusted ), with age of acquisition as the sole predictor (). The All-Model Mix similarly retained only age of acquisition [], explaining of variance (adjusted , ).